OpenStudy (chloed18):

P (x,y) is 6 units from A (0,0) and 9 units from B (9,0). Find x and y.

satellite73 (satellite73):

you have two equations from this by pythagoras \[x^2+y^2=6^2\] which says \((x,y)\) is 6 units from \((0,0)\)

satellite73 (satellite73):

if it is 9 units from \((9,0)\) then \((x-9)^2+y^2=9^2\) now you can solve this \[x^2+y^2=36\\ (x-9)^2+y^2=81\]

OpenStudy (chloed18):

How do you solve it after that?

OpenStudy (chloed18):

@satellite73

satellite73 (satellite73):

idk i didn't do it probably subsitution

satellite73 (satellite73):

for example first one says \(y^2=36-x^2\) so you can substitute that in to the second equation and solve \[(x-9)^2+36-x^2=81\]

OpenStudy (chloed18):

Okay thank you! @satellite73